OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. sin 57° cos 13° - cos 57° sin 13°

OpenStudy (raden):

hint : sin(a)cos(b) - cos(a)sin(b) = sin(a-b)

OpenStudy (anonymous):

57-13=sin(44)? ok thanks!